Required: 1. Complete the following table assuming use of straight-line depreciation. Net cash flow equals the amount of inco2. Complete the following table assuming use of MACRS depreciation. Net cash flow equals the amount of income before deprecia3. Compute the net present value of the investment if straight-line depreciation is used. Use 8% as the discount rate. Chart4. Compute the net present value of the investment if MACRS depreciation is used. Use 8% as the discount rate. Chart Values a(The following information applies to the questions displayed below.) Manning Corporation is considering a new project requir
Answer #1

1.

A B C= A-B D= C*30% E = A - D
Year Income Before depreciation SLM Depreciation Taxable Income Taxes Cashflows
1 67500 9000 58500 17550 49950
2 67500 18000 49500 14850 52650
3 67500 18000 49500 14850 52650
4 67500 18000 49500 14850 52650
5 67500 18000 49500 14850 52650
6 67500 9000 58500 17550 49950

2.

A B C= A-B D= C*30% E = A - D
Year Income Before depreciation MACRS Depreciation Taxable Income Taxes Cashflows
1 67500 18000 49500 14850 52650
2 67500 28800 38700 11610 55890
3 67500 17280 50220 15066 52434
4 67500 10368 57132 17139.6 50360.4
5 67500 10368 57132 17139.6 50360.4
6 67500 5184 62316 18694.8 48805.2

3.NPV (SLM)

Year Cashflow PV Factor @ 8% Discounted Cashflow
1 49950 0.92593 46250.00
2 52650 0.85734 45138.89
3 52650 0.79383 41795.27
4 52650 0.73503 38699.32
5 52650 0.68058 35832.71
6 49950 0.63017 31476.97
PV of cash inflows 239193.16
Less: Cash Outflow 90000.00
NPV 149193.16

4. NPV (MACRS)

Year Cashflow PV Factor @ 8% Discounted Cashflow
1 52650 0.92593 48750.00
2 55890 0.85734 47916.67
3 52434 0.79383 41623.80
4 50360.4 0.73503 37016.40
5 50360.4 0.68058 34274.44
6 48805.2 0.63017 30755.55
PV of cash inflows 240336.86
Less: Cash Outflow 90000.00
NPV 150336.86

Recommendation: Since NPV under MACRS is higher, it is recommended to use MACRS for claiming depreciation.